# Precomputed Dominant Resource Fairness

### v2

@@ 1 Introduction
The rest of this paper is organised as follows: In the next section, we briefly review the literature for related work. In Section 4, we present our model, after reviewing DRF and EDRF in subsections 3.1 and 3.2, respectively, and introducing an analysis method in subsection 3.3. We introduce PDRF in Section 4. In section 5, we present the results of the tests we carried out with PDRF, and departing on the observations gathered by there, in Section 6, we present an heuristic method to improve pure PDRF results. In Section 7 we analyse PDRF performance. Section 8 consists of three subsections, in which we discuss higher order cycles in PDRF (8.1), Pareto Efficiency property (8.2), and weighting policies (8.3), which we believe, are important points to focus on critically. Finally, we conclude in Section 9.
@@ 2 Related Work
3 The Model
We start this section by briefly presenting DRF, and EDRF, and then proceed to introducing PDRF. The notions we introduce in the preceding subsections will help us to construct our own model.
@@ 3.2 Extended Dominant Resource Fairness
3.3 Cyclic Structure of the DRF Main Loop
@@ 6 An Heuristic Method for Allocating Remaining Resources
In addition to the improved accuracy, with the additional heuristic step, the algorithm also reaches Pareto Efficiency. Since overallocations are invariably for the smaller $ds$ values, PDRF makes higher numbers of task allocations with respect to DRF.
@@ 7 Performance of PDRF
As such PDRF offers considerable speed up for large scale applications, such as computing clouds, where the ratio of the resource reserves to the client demands are quite large. Since scheduling is a real time process, in which even small, low level speed ups make an important difference, this point is emphasised even more critically. Cross comparative results from real data may reveal a more accurate measurement, but at this time we cannot provide it, since we do not have access to real data.
@@ 9 Conclusion
One particular contribution of PDRF is that it may be easier to adapt to blockchain context. In fact, it was during our attempt to do so, we came up with the idea, and it lead to the development of a new algorithm for the original context, almost by accident. At that point we turned to a more careful review of the literature, to find EDRF, and the whole endeavour lead to the present study. We are currently working on the final touches of the implementation of the blockchain adaptation, which is a forthcoming study.

### v3

@@ 1 Introduction
The rest of this paper is organised as follows: In the next section, we briefly review the literature for related work. In Section 3, we review DRF and EDRF, and introduce an analysis method in Section 4. We introduce PDRF in Section 5. In section 6, we present the results of the tests we carried out with PDRF, and departing on the observations gathered by there, in Section 7, we present an heuristic method to improve pure PDRF results. In Section 8 we analyse and show the speed-up factor PDRF contributes with respect to the original algorithm. Section 9 consists of three subsections, in which we discuss higher order cycles in PDRF (9.1), Pareto Efficiency property of DRF (9.2), and weighting policies (9.3), which we believe, are important points to focus on critically. Finally, we conclude in Section 10.
@@ 2 Related Work
3 Related Models
In this section we review DRF and EDRF in further detail for their close relevance to our work. The notions introduced in this section will be used to construct our own model in the following sections.
@@ 3.2 Extended Dominant Resource Fairness
4 Cyclic Structure of the DRF Main Loop
@@ 6 Results
Figure 1 shows a sample case for the number of tasks allocated by DRF and PDRF in an overlapping scatterplot. The tests is run with $200$ users for the clarity of the view, from the demand interval 11001-100, with $10$ resources, each resource drawn from 1000001000001100000-1000001 interval.
@@
Figure 1: Scatterplot for number of allocations made by DRF (circles) and PDRF (crosses), with $200$ users and $10$ resources.
@@
Figure 2: Scatterplot for number of allocations made by DRF (circles) and PDRF with heuristic method (crosses), with $200$ users and $10$ resources.
@@ 7 An Heuristic Method for Allocating Remaining Resources
In addition to the improved accuracy, with the additional heuristic step, the algorithm also reaches Pareto Efficiency. Since overallocations are invariably for the smaller $ds$ values, PDRF makes higher numbers of task allocations with respect to DRF. Figure 2 shows the same case in Figure 1 with the updated PDRF allocations after the running of the heuristic method.
@@ 8 Performance of PDRF
As such PDRF offers a significant speed up for large scale applications, such as computing clouds, where the ratio of the resource reserves to the client demands are quite large. Since scheduling is a real time process, in which even small, low level speed ups make an important difference, this point is emphasised even more critically. Cross comparative results constructed with real data may reveal a more accurate measurement, but at this time we cannot provide it, since we do not have access to this data.
@@ 10 Conclusion
One particular contribution of PDRF is that it is easier to adapt to blockchain context. In fact, it was during our attempt to do so, we came up with the idea, and it lead to the development of a new algorithm for the original context, almost by accident. At that point we turned to a more careful review of the literature, to find EDRF, and the whole endeavour lead to the present study. The blockchain adaptation of PDRF is a forthcoming study.